Suppose the first-order maximum for 475 nm blue light being transmitted through a diffraction grating is at an angle of 22°.

Randomized Variables
λ = 475 nm
θ = 22°


How many lines per centimeter are there on this diffraction grating?
1
Expert's answer
2020-12-02T07:14:44-0500

λ = 475 nm "= 475 \times 10^{-9} \;m"

Equation for mth order maximum:

dsinθ = mλ

d – separation between lines on grating

"d = \frac{mλ}{sinθ} \\

d = \frac{1 \times 475 \times 10^{-9}}{0.3746} \\

= 1.268 \times 10^{-6} \;m \\

= 1.268 \times 10^{-4} \;cm"

Number of lines per centimeter (N) of grating:

"N = \frac{1}{d} \\

N = \frac{1}{1.268 \times 10^{-4}} = 7886"

Answer: 7886